# Mapping the transverse spin sum rule in position space

### v1

@@ 4 Conclusions
We then analyzed how the relativistic spatial distributions of transverse OAM, intrinsic spin, and TAM evolve as spin-0 and spin-1/2 targets are boosted along the longitudinal axis. For spin-0 targets, although the spin sum rule is trivially satisfied, it becomes non-trivial at the level of the spatial distribution, and only yields the expected result upon integration. We also discussed the origin and interpretation of each contribution to the transverse OAM distribution in a spin-0 target. For spin-1/2 targets, relativistic intrinsic spin dynamics complicates the internal structure. We examined the transverse OAM, intrinsic spin, and TAM distributions for both unpolarized and transversely polarized spin-1/2 targets. Interestingly, for longitudinally polarized spin-1/2 targets, the transverse AM distributions exactly match those of unpolarized targets. We verified that the transverse spin sum rule is satisfied by integrating over the spatial coordinates the transverse TAM distributions in a transversely polarized spin-1/2 target.

### v2

@@ 4 Conclusions
We then analyzed how the relativistic spatial distributions of transverse OAM, intrinsic spin, and TAM evolve as spin-0 and spin-1/2 targets are boosted along the longitudinal axis. For spin-0 targets, the spin sum rule is trivially satisfied, but the spatial distribution reveals a non-trivial structure that arises purely from Lorentz boost effects. Since this is a spin-independent feature, it also appears in higher-spin targets and plays a key role in understanding how the spatial distribution of angular momentum changes across different frames. Thus, we also discussed the origin and interpretation of each contribution to the transverse OAM distribution in a spin-0 target. For spin-1/2 targets, relativistic intrinsic spin dynamics complicates the internal structure. We examined the transverse OAM, intrinsic spin, and TAM distributions for both unpolarized and transversely polarized spin-1/2 targets. Interestingly, for longitudinally polarized spin-1/2 targets, the transverse AM distributions exactly match those of unpolarized targets. Finally, we verified that the transverse spin sum rule is satisfied by integrating over the spatial coordinates the transverse TAM distributions in a transversely polarized spin-1/2 target.